# Evaluating and improving wave and non-wave stress parametrisations for oceanic flows

### v1

@@ 4 Results and Evaluation
$\text{Stress}(t)=A\cos(\Omega t+\phi)$ (26)
for some determinable values of the amplitude $A$, frequency $\Omega$ and phase $\phi$. To compute these values, we fit the last 200 simulation outputs (or 40 hours) to a sine curve using the Python package pyestimate (Humblet, 2024). To calculate uncertainty, we then recompute these values for the last 100 outputs and the last 101–200 outputs, recording the difference from the initially obtained values.
@@
Finally, for the mixed flow simulations, we assume the stress was of the form (cf. (21))
@@
with both a mean component $F_{m}$ and a periodic component $F_{p}\cos(\Omega t+\phi)$. To compute $F_{m}$ and $F_{p}$ we again use the pyestimate package to fit the last 200 simulation outputs to a shifted sine curve of the form (27). The uncertainty in $F_{m}$ and $F_{p}$ is obtained by recomputing these values using the last 100 outputs and 101–200 outputs and recording any difference.
@@ 4.1 Tidal flow in two dimensions
We first consider tidal flow over an isolated hill in two dimensions. For each simulation, the computed stress is compared to the parametrisations in Section 22.1. Figure 2 shows this comparison for the parametrisation $F_{\text{SAH2d}}$ (2.1.2) when the parameters $U_{\text{tidal}}$, $N$, $f$, $W$, $H$ and $h_{0}$ are varied. For most of these simulations, the flow is forced with the semi-diurnal $M_{2}$ constituent, with $\omega=1.4\times 10^{-4}\ $\mathrm{s}$^{-1}$. The only exception is the bottom-trapped $(|f|>\omega)$ tests, whereby we use the diurnal $K_{1}$ constituent ($\omega=7.3\times 10^{-5}\ $\mathrm{s}^{-1}$$). This is because bottom-trapped effects are more common for diurnal constituents than semi-diurnal constituents (since the frequency $\omega$ is smaller).
@@
In general, the parametrisation $F_{\text{SAH2d}}$ performs well compared to the simulations. However, in most cases the theoretical prediction slightly underestimates the stress. This is to be expected as the parametrisations from JSL2001 and SAH2020 are derived in an idealised regime whereby the topographical height and flow speed are assumed to be small. There are thus other non-linear effects that are not accounted for which could contribute to the drag. For instance, calculations of Balmforth et al. (2002) show that the energy flux should increase by around 14% above the linear theory for isolated Gaussians as the steepness parameter
$s=\frac{Nkh_{0}}{\sqrt{\omega^{2}-f^{2}}}$ (28)
goes from 0 to 1, which we expect should also translate to form drag. Note that with the default parameters used here, $s\approx 0.6$ for the dominant wavenumber.
@@
Looking at Figures 2a and 2b, we see that the predicted linear scaling by $U_{\text{tidal}}$ and $N$ matches the simulations almost precisely. The same is true for the $\sqrt{\omega^{2}-f^{2}}/\omega$ scaling for $f$ (Figure 2c), highlighting the importance of accounting for latitude in tidal stress, which was neglected in the JSL2001 parametrisation.
@@
However, for the simulations varying $W$, $H$ and $h_{0}$ (Figures 2d–2h), there are some larger discrepancies between the computed stress and $F_{\text{SAH2d}}$. Firstly for the tests varying $W$ we see a small increase in the drag as $W$ increases. Note that a priori we do expect some dependence of the stress on $W$ despite it not appearing in $F_{\text{SAH2d}}$. This is because a wider hill gives more area for waves reflected off the surface to interact with topography. However, it is not clear whether this should always correspond to an increase in stress as observed in the simulations.
@@
Next, for the tests varying $H$ and $h_{0}$ we find that, provided $|f|<\omega$, the flow in the vicinity of the hill speeds up as it moves over the narrow channel at the top of the hill. Thus, to better model the stress for the $H$ and $h_{0}$ tests we use the velocity scaling
@@
$\widetilde{U}_{m}(t)=\begin{cases}\frac{H}{H-h_{0}}U(t),&h_{0}\leq H/2\\ 2\>U(t),&h_{0}>H/2\end{cases}$ (30)
@@
yields better results for large values of $h_{0}$. This is shown in Figures 2e–2g, whereby incorporating $\widetilde{U}_{m}(t)$ into the $F_{\text{SAH2d}}$ parametrisation gives a stress profile that closely matches the simulation data. The only exception is when the hill crest is close to the surface ($h_{0}\approx H$) and the stress plateaus (Figure 2g).
@@
For bottom-trapped ($|f|\geq\omega$) flows we found that the $F_{\text{SAH2d}}$ parametrisation remains relatively accurate as $h_{0}$ increases (Figure 2h). This is expected as such flows have a different velocity profile compared to the $|f|<\omega$ case. Namely, when $|f|\geq\omega$, the flow still speeds up at the top of the hill but is slowed down or “trapped” at the base of the hill. Thus in this case, it is not accurate to use the $\widetilde{U}_{m}$ scaling, which assumes the current speeds up along the entire length of the hill.
@@
Finally, we analyse the frequency and phase of the simulated stress and how it compared to the parametrisations. For the tests varying $f$, this is shown in Figure 3. Here, the $F_{\text{SAH2d}}$ parametrisation suggests that the frequency of the stress should be equal to $\omega$. Then, in the absence of wave reflections, the stress should be in phase with $U(t)$ when $|f|<\omega$ and then $90^{\circ}$ (or 1/4 of a tidal period) out of phase with $U(t)$ when $|f|>\omega$. Note that we give the phase in hours, obtained by multiplying $\phi$ (from (26)) by $\omega/2\pi$. Most of our simulations accurately reflect the predicted values from $F_{\text{SAH2d}}$, usually giving a stress profile with the predicted frequency (Figure 3a) and within half an hour of the predicted phase (Figure 3b). This provides further support for using the SAH2020 parametrisation over the JSL2001 parametrisation, which does not account for the phase shift when $|f|>\omega$.
Figure 2: The amplitude of the oscillatory stress given by $F_{\text{SAH2d}}$ and the simulation data for two-dimensional tidal flow. The default parameters for each test are as in Table 1 with $\omega=10^{-4}\ $\mathrm{s}^{-1}$$ ($M_{2}$ tide) at a latitude of $20^{\circ}S$ ($f=5\cdot 10^{-5}\ $\mathrm{s}^{-1}$$). The only exception was the bottom-trapped tests (Figure (h)) which used $\omega=7.3\cdot 10^{-5}\ $\mathrm{s}^{-1}$$ ($K_{1}$ tide) at a latitude of $90^{\circ}S$ ($f=1.46\cdot 10^{-4}\ $\mathrm{s}^{-1}$$). In all tests, the uncertainty in the computed stress was less than 3% so the (small) error bars have been omitted.
Figure 3: The fitted frequency and phase of the stress for our two-dimensional tidal flow simulations varying latitude. Here, the phase $\phi^{\prime}=\phi\times(\omega/2\pi)$ is given in hours as opposed to radians (cf. (26)). In each test, the uncertainty was less than $10^{-6}$ rad/s in the frequency, and less than $0.05$ hours in the phase so the (small) error bars have been omitted.
@@ 4.2 Tidal flow in three dimensions
Our results for tidal flow in three dimensions are analysed analogously to our results in two dimensions. Note that in what follows, we compute the $x$-component of the stress amplitude, frequency and phase. Figure 4 shows a comparison between $F_{\text{SAH3d}}$ and the simulation data when the parameters $U_{\text{tidal}}$, $W$, $f$ and $h_{0}$ were varied. Plots of the stress frequency and phase are also included for tests that vary the latitude (or equivalently $f$).
Similar to the two-dimensional case, the simulations varying $U_{\text{tidal}}$, $W$ and $f$ (Figures 4a–4c) qualitatively reflect the scaling predicted by the $F_{\text{SAH3d}}$ parametrisation. However, for tests varying $f$ there is a notably larger quantitative deviation between the simulated stress and that predicted by $F_{\text{SAH3d}}$. This is particularly true near the critical latitude where $|f|=\omega=1.4\cdot 10^{-4}\ $\mathrm{s}^{-1}$$. To account for this, we note that the $F_{\text{SAH3d}}$ parametrisation is derived assuming a small excursion parameter $\epsilon=k\overline{u}/\omega$. As a result of this assumption, the temporal and spatial terms of the governing equations cancel when $|f|=\omega$. Therefore, the fact that the simulated stress is non-zero at the critical latitude represents the limitation of such an assumption, and our results indicate that this effect may be more impactful in three dimensions.
We now consider tests varying $h_{0}$. In the bottom-trapped ($|f|>\omega$) case, the simulated data matches the scaling predicted by the $F_{\text{SAH3d}}$ parametrisation (Figure 4d). However, for our choice of parameters, the simulated stress underestimates $F_{\text{SAH3d}}$ more than in the two-dimensional case. In particular, $F_{\text{SAH3d}}$ overestimated the simulation data by an average of $146\%$ in three dimensions, compared to $118\%$ in two-dimensions. In the case where $|f|<\omega$ there is then a qualitative difference between the computed stress profile and that predicted by $F_{\text{SAH3d}}$, depending on whether the hill width $W$ is small or large (Figures 4f and 4g). To account for this difference, we let $\widetilde{F}_{\text{SAH3d}}$ be equal to $F_{\text{SAH3d}}$ but with the velocity scaling $\widetilde{U}_{m}$ (Equation (30)) discussed in the two-dimensional setting. For large $W$ (Figure 4f), the scaling of the simulated stress is the same as in the two-dimensional case. Namely, it matches the stress profile predicted by $\widetilde{F}_{\text{SAH3d}}$ until $h_{0}\approx H$. However, for our default width $W=5000\ m$, the simulated stress only matches $\widetilde{F}_{\text{SAH3d}}$ for small–medium values of $h_{0}$ (Figure 4e). Then, for $h_{0}\gtrsim 500\ m$, the stress is much closer to the unscaled $F_{\text{SAH3d}}$ parametrisation. This is not entirely unexpected as provided $W$ is not too large, we expect some of the current to flow around the hill, rather than be forced over the top as in the two-dimensional case.
Finally, we remark that, as with the two-dimensional case, the frequency and phase accurately matches the values predicted by the SAH2020 theory (Figures 4g and 4h). However, the most notable difference in three dimensions is a smoother transition from the stress being in phase when $|f|<\omega$ to being $90^{\circ}$ out of phase when $|f|>\omega$. Nevertheless, this effect was only clear when $|f|$ was very close to $\omega$ so that the phase predicted in $F_{\text{SAH3d}}$ is still accurate for almost all values of $f$.
@@
Figure 4: The amplitude of the oscillatory stress given by $F_{\text{SAH3d}}$ and the simulation data for three-dimensional tidal flow (panels (a)–(f)), and the frequency and phase of the stress as the latitude is varied (panels (g) and (h)). The default values for each test were the same as the two-dimensional case (see Figure 2 and Table 1). In panels (a)–(g), the uncertainty in the computed values was less than 10% so the (small) error bars have been omitted. In Figure (h), error bars have only been included in the single case where the uncertainty in the phase was greater than 0.1 hours.

### v3

@@ 4 Results and Evaluation
$\text{Stress}(t)=A\cos(\Omega t+\phi),$ (30)
for some determinable values of the amplitude $A$, frequency $\Omega$ and phase $\phi$. To compute these values, we first extract the last 200 values (corresponding to 40 hours) of the stress output by the model, computed as per (1) or (28), depending on the dimension under consideration. We then fit these 200 values to a cosine curve. To calculate uncertainty, we recompute these values for the last 100 values and the last 101–200 values, recording the difference from the initially obtained values.
@@
Finally, for the mixed flow simulations, we assume the stress is of the form (cf.  (21))
@@
with both a mean component $F_{m}$ and a periodic component $F_{p}\cos(\Omega t+\phi)$. To compute $F_{m}$ and $F_{p}$ we fit the last 200 output values of the stress to a shifted cosine curve of the form (31). The uncertainty in $F_{m}$ and $F_{p}$ is obtained by recomputing for the last 100 values and 101–200 values and recording any difference.
Table 1: The default parameter values used for the tests with a purely tidal or steady flow in either two dimensions (2d) or three dimensions (3d). Here NN is the buoyancy frequency, $h_{0}$ is the hill height, $W$ is the hill width, $H$ is the ocean depth, $f$ is the Coriolis parameter, $\omega$ is the tidal frequency, $U_{\text{mean}}$ is the mean flow speed and $U_{\text{tidal}}$ is the amplitude of the periodic component of the flow. For the tidal flow tests, the choice of $\omega$ and $f$ (and thus latitude) will be made clear in the text. The parameters for the steady flow tests also vary depending on whether we are interested in a low or high Froude number $Fr=Nh_{0}/U_{\text{mean}}$ regime.
Tidal Steady $(Fr\lesssim 1)$ Steady ($Fr\gg 1$)
NN $(\mathrm{s}^{-1})$ $4\times 10^{-3}$ $2\times 10^{-3}$ $5\times 10^{-3}$
$h_{0}$ $(\mathrm{m})$ $100$ 2020 (2d) or $30$ (3d) $500$
$W$ $(\mathrm{m})$ $5000$ $5000$ $5000$
$H$ $(\mathrm{m})$ $1500$ $1500$ (2d) or $1250$ (3d) $1500$
$f$ $(\mathrm{s}^{-1})$ $5\times 10^{-5}$ or $1.46\times 10^{-4}$ $2.53\times 10^{-5}$ 0
Latitude $(^{\circ}\textrm{S})$ 20 or 90 10 0
$\omega$ $(\mathrm{s}^{-1})$ $1.4\times 10^{-4}$ or $7.3\times 10^{-5}$ - -
[…]
$Fr=Nh_{0}/U_{\text{mean}}$ - $0.2$ (2d) or $0.3$ (3d) $25$
[…]
@@ 4.1 Tidal flow in two dimensions
Figure 2: The amplitude of the oscillatory stress given by $F_{\text{SAH2d}}$ and the simulation data for two-dimensional tidal flow. In Figures (e)–(g) a refinement of the $F_{\text{SAH2d}}$ parametrisation with a scaled velocity (see Equation (33)) is also shown in red. The default parameters for each test are as in Table 1 with $\omega=10^{-4}\ $\mathrm{s}^{-1}$\$ ($M_{2}$ tide) at a latitude of $20^{\circ}S$ ($f=5\times 10^{-5}\ $\mathrm{s}^{-1}$\$). The only exception is the bottom-trapped tests in Figure (h) which use $\omega=7.3\times 10^{-5}\ $\mathrm{s}^{-1}$\$ ($K_{1}$ tide) at a latitude of $90^{\circ}S$ ($f=1.46\times 10^{-4}\ $\mathrm{s}^{-1}$\$). In all tests, the uncertainty in the computed stress was less than 3% so the small error bars have been omitted.
Figure 3: The fitted phase of the stress for our two-dimensional tidal flow simulations varying latitude. Here, the phase $\phi^{\prime}=\phi\times(\omega/2\pi)$ is given in hours as opposed to radians (cf. (30)), and the critical latitude ($\approx 74.5^{\circ}S$) is displayed in red. In each test, the uncertainty was less than $0.05$ hours so the small error bars have been omitted.
We first consider tidal flow over an isolated hill in two dimensions. For each simulation, the computed stress is compared to the parametrisations in Section 22.1. Figure 2 shows this comparison for the parametrisation $F_{\text{SAH2d}}$ (see (2.1.2)) when the parameters $U_{\text{tidal}}$, $N$, $f$, $W$, $H$ and $h_{0}$ are varied. Although not explicitly shown in Figure 2, the simpler JSL2001 parametrisation (Equation (3)) predicts the same scaling as $F_{\text{SAH2d}}$ with respect to the stratification $N$, hill height $h_{0}$ and tidal speed $U_{\text{tidal}}$. Thus our analysis for these variables also holds true for the JSL2001 theory.
For most of the simulations, the flow is forced with the semi-diurnal $M_{2}$ constituent, with $\omega=1.4\times 10^{-4}\ $\mathrm{s}$^{-1}$. The only exception is for the bottom-trapped $(|f|>\omega)$ tests displayed in Figure 2h, whereby we use the diurnal $K_{1}$ constituent ($\omega=7.3\times 10^{-5}\ $\mathrm{s}^{-1}$\$). This is because on Earth, bottom-trapped effects are more common for diurnal constituents than semi-diurnal constituents. In particular, for a diurnal constituent the frequency $\omega$ is smaller and thus $|f|>\omega$ for a larger range of latitudes. However, note that we also performed a small number of tests with $|f|>\omega$ for the $M_{2}$ tide. This appears as part of our tests varying $f$ (Figure 2c), in the case when the latitude is greater than $75^{\circ}$.
@@
In general, the parametrisation $F_{\text{SAH2d}}$ performs well compared to the simulations. However, in most cases the theoretical prediction slightly underestimates the stress. This is to be expected as the parametrisations from JSL2001 and SAH2020 are derived in an idealised regime whereby the topographic height and flow speed are assumed to be small. There are thus other non-linear effects that are not accounted for which could contribute to the drag. For instance, calculations of Balmforth et al. (2002) show that the energy flux should increase by around 14% above the linear theory for isolated Gaussian hills as the steepness parameter (defined in (29)) goes from 0 to 1, which we expect should also translate to form stress.
@@
Looking at Figures 2a and 2b, we see that the predicted linear scaling by the tidal speed $U_{\text{tidal}}$ and the buoyancy frequency $N$ matches the simulations almost precisely. The same is true for the $\sqrt{\omega^{2}-f^{2}}/\omega$ scaling for the Coriolis parameter $f$ (Figure 2c), highlighting the importance of accounting for latitude in tidal stress, which is neglected in the JSL2001 parametrisation.
@@
However, for the simulations varying the hill width $W$, ocean depth $H$ and hill height $h_{0}$ (Figures 2d–2h), there are some larger discrepancies between the computed stress and $F_{\text{SAH2d}}$. Firstly for the tests varying $W$ we see a small increase in the drag as $W$ increases. Note that a priori we do expect some dependence of the stress on $W$ despite it not appearing in $F_{\text{SAH2d}}$. This is because a wider hill gives more area for waves reflected off the surface to interact with topography, as discussed around (7). Additionally, we also note that the slope criticality (see (29)) directly depends on the wavenumber $k$ and thus $W$.
@@
Next, for the tests varying the ocean depth $H$ and hill height $h_{0}$ (Figures 2e–g) we find that, provided $|f|<\omega$, the flow in the vicinity of the hill speeds up as it moves over the narrow channel at the top of the hill. As a result, one finds that $F_{\text{SAH2d}}$ underestimates the stress particularly when $h_{0}$ is large or $H$ is small. Thus, to better model the stress for the $H$ and $h_{0}$ tests we also consider a refinement of the $F_{\text{SAH2d}}$ parametrisation with a scaled velocity
@@
$\widetilde{U}_{m}(t)=\begin{cases}\dfrac{H}{H-h_{0}}U(t),&h_{0}\leq H/2\\ 2\>U(t),&h_{0}>H/2\end{cases}$ (33)
@@
yields better results for large values of $h_{0}$. In particular, for our tests varying $h_{0}$ in Figures 2f and 2g, we see that incorporating $\widetilde{U}_{m}(t)$ into the $F_{\text{SAH2d}}$ parametrisation gives a stress profile that is much closer to the simulation data. The only exception is when the hill crest is close to the surface ($h_{0}\approx H$) and the stress plateaus (Figure 2g). We remark that for our tests, there is a discontinuity in the derivative (with respect to $h_{0}$) of $\widetilde{U}_{m}(t)$ at $h_{0}=750\ $\mathrm{m}$\$. However, this is not shown explicitly in Figure 2 since the discontinuity is at the point which separates Figures 2f and 2g. For our tests varying $H$, a similar improvement upon using $\widetilde{U}_{m}(t)$ can then be seen in Figure 2e. However, even with the $\widetilde{U}_{m}(t)$ scaling, $F_{\text{SAH2d}}$ still underestimates the model output as $H$ decreases. This is not unexpected as wave reflections become more important in this limit (see discussion around (7)).
@@
We now consider the results for bottom-trapped flows with $|f|\geq\omega$ (Figure 2h), which were performed for the K1 tidal constituent at a latitude of $90^{\circ}$. For these tests, in contrast to the $|f|<\omega$ case, we found that the $F_{\text{SAH2d}}$ parametrisation remains relatively accurate as $h_{0}$ increases (Figure 2h). This is expected as such flows have a different velocity profile compared to the $|f|<\omega$ case. Namely, when $|f|\geq\omega$, the flow still speeds up at the top of the hill but is slowed down or “trapped” at the base of the hill. Thus in this case, it is not accurate to use the $\widetilde{U}_{m}$ scaling, which assumes the current speeds up along the entire length of the hill.
@@
Finally, as defined in (30), it remains to analyse the frequency $\Omega$ and phase $\phi$ of the simulated stress and how it compares to the parametrisations. Here, we first remark that the fitted frequency from our simulation data closely matches the tidal frequency: in each case the two quantities are within 1% of each other. However, the phase data is more interesting and for the tests varying $f$, this is shown in Figure 3. Here, the $F_{\text{SAH2d}}$ parametrisation suggests that in the absence of wave reflections, the stress should be in phase with $U(t)$ when $|f|<\omega$ and then $90^{\circ}$ (or 1/4 of a tidal period) out of phase with $U(t)$ when $|f|>\omega$. Note that we give the phase in hours, obtained by multiplying $\phi$ by $\omega/2\pi$. Most of our simulations accurately reflect this prediction, yielding a phase within half an hour of that predicted by $F_{\text{SAH2d}}$. The only exception is when the latitude is very close to the critical value where $|f|=\omega=1.4\times 10^{-4}s$ or approximately $74.5^{\circ}S$. In this exceptional case, the phase is near the half-way point between the predicted subcritical ($|f|<\omega$) and supercritical ($|f|>\omega$) phase. This behaviour is exactly what one would expect from Fourier analysis. However, there would be little practical value in adding this exceptional case to the $F_{\text{SAH2d}}$ parametrisation. Ultimately though, our results for the phase further support including an $f$ dependency in stress parametrisations, which is not present in the JSL2001 parametrisation.
@@ 4.2 Tidal flow in three dimensions
Figure 4: The amplitude of the oscillatory stress given by $F_{\text{SAH3d}}$ and the simulation data for three-dimensional tidal flow (panels (a)–(f)), and the phase of the stress as the latitude is varied (panel (g)). The default values for each test are the same as the two-dimensional case (see Figure 2 and Table 1). In panels (a)–(f), the uncertainty in the computed values was less than 10% so the small error bars have been omitted. In Figure (g), error bars have only been included in the single case where the uncertainty in the phase was greater than 0.1 hours.
Our results for tidal flow in three dimensions are analysed analogously to our results in two dimensions. Note that in what follows, we compute the $x$-component of the stress amplitude, frequency and phase. Figure 4 shows a comparison between $F_{\text{SAH3d}}$ (Equation (2.1.2)) and the simulation data as the tidal speed $U_{\text{tidal}}$, hill width $W$, Coriolis parameter $f$ and hill height $h_{0}$ are varied. Plots of the phase are also included for tests that vary the latitude (or equivalently $f$).
Similar to the two-dimensional case, the simulations varying $U_{\text{tidal}}$, $W$ and $f$ (Figures 4a–4c) qualitatively reflect the scaling predicted by the $F_{\text{SAH3d}}$ parametrisation.
For the tests varying $f$, the $F_{\text{SAH3d}}$ parametrisation well approximates the model output for near-zero $|f|$. However, there is an almost linearly increasing error as $f$ increases. As this effect did not occur in two-dimensions, there must be something specific to three-dimensions causing this discrepancy. As $|f|$ increases, there is increasing flow in the $y$ direction which in three-dimensions (unlike two) is permitted to vary in $y$, and we speculate that this additional degree of freedom is likely responsible for the growing error, although the details remain uncertain.
We now consider tests varying $h_{0}$. In the bottom-trapped ($|f|>\omega$) case, the simulated data matches the scaling predicted by the $F_{\text{SAH3d}}$ parametrisation (Figure 4d). However, for our choice of parameters, $F_{\text{SAH3d}}$ overestimates the simulated stress by more than in the two-dimensional case. In particular, over all the tests we performed, $F_{\text{SAH3d}}$ was on average $46\%$ larger than the simulated stress in three dimensions, compared to $18\%$ for $F_{\text{SAH2d}}$ in two-dimensions. In the case where $|f|<\omega$ there is then a qualitative difference between the computed stress profile and that predicted by $F_{\text{SAH3d}}$, depending on whether the hill width $W$ is small or large (Figures 4e and 4f). To account for this difference, we let $\widetilde{F}_{\text{SAH3d}}$ be equal to $F_{\text{SAH3d}}$ but with the velocity scaling $\widetilde{U}_{m}$ (Equation (33)) discussed in the two-dimensional setting. For large $W$ (Figure 4f), the scaling of the stress is the same as in the two-dimensional case. Namely, $\widetilde{F}_{\text{SAH3d}}$ matches the stress profile until $h_{0}\approx H$. However, for our default width W=5000mW=5000\ m, $\widetilde{F}_{\text{SAH3d}}$ only matches the simulated stress for small–medium values of $h_{0}$ (Figure 4e). Then, for $h_{0}\gtrsim 500\ m$, the stress is much closer to the unscaled $F_{\text{SAH3d}}$ parametrisation. Since $\widetilde{F}_{\text{SAH3d}}$ is only slightly larger than $F_{\text{SAH3d}}$ for small hill heights, it thus seems reasonable to just apply the unscaled parametrisation $F_{\text{SAH3d}}$ for all values of $h_{0}$ in this “small WW” regime. We remark that this regime shift for smaller $W$ is not entirely unexpected, as provided $W$ is not too large, we expect some of the current to flow around the hill, rather than be forced over the top as in the two-dimensional case. Future work should identify more precisely at what values of $W$ this regime shift occurs.
Finally, we remark that, as with the two-dimensional case, the frequency and phase accurately matches the values predicted by the SAH2020 theory. However, the most notable difference in three dimensions is a smoother transition from the stress being in phase when $|f|<\omega$ to being $90^{\circ}$ out of phase when $|f|>\omega$ (Figure 4h). Nevertheless, this effect was only clear when $|f|$ was very close to $\omega$ so that the phase predicted in $F_{\text{SAH3d}}$ is still accurate for almost all values of $f$.